OpenStudy (anonymous):

Which equation is in standard form and represents a line with slope –1 through the point (–3, 5)?

OpenStudy (anonymous):

Hello! Are you familiar with the point-slope equation?

OpenStudy (anonymous):

Not Really... Im Just Learning it... Can Yu Help?

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (anonymous):

x1 and y1 are the coordinates of the point

OpenStudy (anonymous):

m is the slope

OpenStudy (anonymous):

How can we use this?

OpenStudy (anonymous):

So replace m with -1?

OpenStudy (anonymous):

Correct! And what about x1 and y1?

OpenStudy (anonymous):

y-5 = m(x-3) ?

OpenStudy (anonymous):

the left side is correct!

OpenStudy (anonymous):

Let's take a look at the right side: First, we can put -1 instead of m, right?

OpenStudy (anonymous):

and what about x1?

OpenStudy (anonymous):

Which is the value of x1?

OpenStudy (anonymous):

so y-5 = -1(x-3) ?

OpenStudy (anonymous):

That's better, but we still have to take a look at x1 Which is the value of x1?

OpenStudy (anonymous):

since x1=-3 we will get this:

OpenStudy (anonymous):

y-5 = -1(x+3)

OpenStudy (anonymous):

And then we have to write it in its standard form

OpenStudy (texaschic101):

I do not mean to interrupt, but if you are having a hard time with the point slope form, you can solve this by just using y = mx + b form. Just a thought.